UPDATE - 10:21 A.M. Service to the Neon cluster has been restored and jobs are running normally. The datacenter operations team has confirmed a brief power outage earlier this morning that caused the issue.
Thank you for your patience.
This morning the HPC team identified problems with the Neon cluster. At this time a large number of nodes are offline and we are working to fix the problem. At this time queues are disabled while we restore service.
Apologies for the inconvenience,